OpenStudy (anonymous):

1. Factor and Solve x^2 – 15x + 50=0

OpenStudy (jdoe0001):

hint: \(\begin{array}{cccllll} x^2 &- 15x &+ 50&=0\\ &\uparrow &\uparrow \\ &-10-5&-10\cdot -5 \end{array} \)

OpenStudy (anonymous):

?

OpenStudy (jdoe0001):

well... have you covered factoring quadratics yet?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

well... then thereins lies the issue you may want to cover that first then, check your book on factoring quadratics

OpenStudy (anonymous):

I did. my teachers are MAKING me cover this first

OpenStudy (jdoe0001):

well.... then notice above -10 -5 = -15 -10 * -5 = +50 you could take it from there then :)

OpenStudy (anonymous):

I have to complete these 5 questions first 1. Factor and Solve x^2 – 15x + 50=0 2. What is the Y intercept of x^2 – 15x + 50=0 3. X^2 + 6x + 12 = (x+3) True False 4. Which of the following would be the BEST problem to use the complete the square method? X ^2 + 9x + 20 x^2 - 49 x^2 -3x + 31 x^2 + 8x – 11 5. Which of the following can be factored? (Points : 2) x^2 + 10x + 20 x^2 + 10x + 21 x^2 + 10x + 30 A, B, and C can all be factored

OpenStudy (jdoe0001):

well, take them one at a time if your teacher is making you do this, and you've covered the factoring then he/she is correct, because the 1st one is factoring.... so should be simple enough then